# Think How to Think: Mitigating Overthinking with Autonomous Difficulty Cognition in Large Reasoning Models

### v2

@@ Abstract
Recent Large Reasoning Models (LRMs) excel at complex reasoning tasks but often suffer from overthinking, generating overly long and redundant reasoning trajectories. To explore its essence, our empirical analysis reveals that LRMs are primarily limited to recognizing task properties (i.e., difficulty levels) like humans before solving the problem, leading to a one-size-fits-all reasoning process. Inspired by this, a pressing and natural question emerges: Can we explicitly bootstrap such ability to alleviate overthinking in LRMs? In this paper, we propose Think-How-to-Think (TH2T), a novel two-stage fine-tuning strategy that progressively inspires LRMs’ difficulty cognition and redundancy cognition of LRMs. Specifically, we first inject difficulty hypnosis into output prefixes to guide the model toward adaptive reasoning depth, trained on a hybrid dataset mixing short and long reasoning paths. Then, we incorporate redundancy hypnosis, which supervises the intermediate reasoning steps to identify and eliminate unnecessary reasoning patterns. Experiments on 7B/14B/32B models demonstrate that TH2T significantly reduces inference costs by over 70% on easy tasks and 40% on hard tasks while maintaining performance stability. The resulting outputs exhibit clear signs of difficulty-aware capabilities and reduced redundancy (e.g., reflection and looping).
Introduction
@@
The emergence of reasoning-focused Large Language Models, also known as Large Reasoning Models (LRMs), has led to substantial improvements in problem-solving capabilities (Sun et al. 2024b; Yu et al. 2024; Team et al. 2025), particularly through the use of elaborate and sequential reasoning enabled by test-time scaling (Snell et al. 2025). In addition to OpenAI-O1 (OpenAI et al. 2024a), other models such as the open-source Deepseek-R1 (Guo et al. 2025) and QwQ (Team 2025) leverage long Chain-of-Thought (Wei et al. 2022; Wang et al. 2023) and adopt reinforcement learning (RL) to approach complex tasks step-by-step, enabling a human-like thinking pattern with reflection, backtracking, and self-validation (Guo et al. 2025; Xie et al. 2025; Liu et al. 2025). This reasoning paradigm demonstrably enhances the models’ problem-solving capabilities and yields promising results. However, lengthy responses introduce the problem of “overthinking” (Team et al. 2025; Chen et al. 2025), wherein model performance initially improves with extended reasoning chains but deteriorates beyond a certain length. For instance, even for trivial queries such as “What is 1 plus 1?”, current models often misjudge the difficulty and produce unnecessarily long outputs—sometimes spanning hundreds of tokens. This tendency toward redundant reasoning not only increases computational cost and latency but also degrades user experience. These observations prompt a critical question: do LRMs possess the capacity to recognize and adapt to varying levels of problem difficulty autonomously?
@@
To answer this question, we start from the perspective of cognitive science, specifically the dual-process theory of human reasoning. This theory posits that human cognition operates via two distinct but complementary systems: System 1, which is fast, concise, and intuitive, and System 2, which is slow, deliberate, and analytical (Kahneman 2011; Evans and Stanovich 2013). Before engaging in problem-solving, humans often perform an initial assessment of a task’s difficulty, allowing them to allocate cognitive resources appropriately between these two systems (Efklides 2011; Flavell 1979). For straightforward tasks, reliance on System 1 suffices, whereas complex problems necessitate the engagement of System 2. This metacognitive ability to evaluate task difficulty and adapt reasoning strategies accordingly, however, is lacking in current reasoning models. As shown in Fig. 2, DeepSeek-R1-Distill-Qwen2.5-7B frequently conflates easy datasets (e.g., GSM8K (Cobbe et al. 2021)) and hard datasets (e.g., AIME2024), treating them with a uniform level of reasoning complexity. These limitations highlight the need for LRMs to develop autonomous difficulty cognition—to “think how to think”.
Current research primarily emphasizes unified CoT compression based on length, yet length is not an accurate proxy for difficulty. In this paper, we propose Think-How-to-Think (TH2T), a novel two-stage fine-tuning strategy that progressively inspires LRMs difficulty cognition and redundancy cognition. In contrast to prompt-based input guidance for instruction adherence, our approach emphasizes output-level interventions, directly intervening inside the reasoning trajectory to establish a self-hypnosis mechanism. Specifically, in the first stage, to encourage the model to develop native difficulty cognition autonomously, we design difficulty-awareness hypnosis and inject it into response prefixes to establish cognitive hints. Combined with a heterogeneous dataset, the trained model enhances its sensitivity to task difficulty, thereby reducing overthinking in simpler tasks through structured cognitive self-guidance.
In the second stage, we conduct fine-grained in-progress redundancy management, enhancing LRM’s redundancy cognition by targeting redundant yet logically complete inference structures (e.g., reflections or double-checks) and terminal reasoning loops commonly observed in complex task generations. The identified structures are injected with redundancy hypnosis while truncating subsequent superfluous content. Overall, our contributions can be outlined as follows:
@@
We observe that current LRMs fall short of perceiving task complexity. To mitigate overthinking, we propose enhancing LRMs with the ability to cognize and adapt to varying levels of problem difficulty autonomously.
@@
To align with human System 1 and System 2 thinking patterns, we propose a novel two-stage fine-tuning strategy that progressively inspires LRMs’ difficulty and redundancy cognition with a constructed heterogeneous dataset with inserted self-hypnosis.
@@
Experiments demonstrate over 70% length reduction on easy tasks and 40% on hard tasks with maintained performance stability, exhibiting clear difficulty-aware capabilities and reduced redundancy.
Literature Review
@@
The advent of LRM (OpenAI et al. 2024a; Guo et al. 2025; Team 2025) enables breakthroughs in problem solving (Sun et al. 2024b; Yu et al. 2024) with Chain-of-Thought (Wei et al. 2022; Wang et al. 2023) and test-time scaling (Snell et al. 2025). Nevertheless, this introduces the challenge of overthinking (Chen et al. 2025; Team et al. 2025). Works of efficient reasoning (Sui et al. 2025) as summarized:
Model-based methods.
Reinforcement Learning motivates some existing efficient reasoning works (Arora and Zanette 2025; Team et al. 2025; Aggarwal and Welleck 2025) to incorporate length penalties. O1-pruner (Luo et al. 2025b) introduces the Length-Harmonizing Reward, integrated with a PPO-style loss. DAST (Shen et al. 2025b) constructs a length-preference dataset and employs SimPO (Meng and Xia 2024). AdaR1 (Luo et al. 2025a) constructs a hybrid reasoning model by merging long and short CoT models and is optimized with DPO (Rafailov et al. 2023). Apart from RL, which is vulnerable and expensive, fine-tuning is another effective post-training strategy (Yang et al. 2025b). CoT-Value (Ma et al. 2025b) regulates reasoning length by a controllable direction in parameter space, enabling dynamic length control. C3oT (Kang et al. 2025) employs GPT-4 as a compressor while retaining key information. TokenSkip (Xia et al. 2025) prunes less important tokens and fine-tune the model to skip the redundant tokens.
Output-based methods.
Implicit-based methods (Deng, Choi, and Shieber 2024; Hao et al. 2024; Saunshi et al. 2025; Cheng and Durme 2024; Shen et al. 2025a) propose to internalize reasoning steps within the latent representation, transitioning reasoning from language space to hidden space. (Saunshi et al. 2025) simulates multiple steps of CoT through a looped transformer layer, increasing the depth of reasoning without additional parameters. However, these implicit techniques introduce performance decline and affect interpretability. There are also dynamic criteria. For example, (Sun et al. 2024a; Xie et al. 2023; Liao et al. 2025) optimize Best-of-N (BoN) decoding with speculation and rejection. (Li et al. 2024; Yang et al. 2025a; Wang et al. 2025) explores possible early exit point. NoThinking (Ma et al. 2025a) prompts reasoning models to directly output final answers without thinking, but heavily relies on sampling, which adversely increases overall overhead.
Prompt-based methods.
@@
Some works (Aytes, Baek, and Hwang 2025; Ding et al. 2024; Ning et al. 2024; Han et al. 2025; Xu et al. 2025; Lee, Che, and Peng 2025) instruct LRMs by various prompts to respond concisely. Specifically, (Renze and Guven 2024) explicitly prompts LLMs to perform concise step-by-step reasoning. SoT (Ning et al. 2024) combines reasoning paradigms with linguistic constraints to minimize token usage. Token-Budget (Han et al. 2025) establishes a token budget estimated by LLM itself, and incorporates this estimation as a token number constraint. (Ge et al. 2025) regulates thinking token distribution by few-shot in-context learning. Nevertheless, prompting relies on precise instruction following and robust task cognition, which will be proven unreliable later.
Overall, most existing works treat efficient reasoning as a unified length compression paradigm. There is a necessity that has been widely overlooked that LRM may need the capacity to autonomously recognize and adapt to varying levels of problem difficulty, which aligns with System 1 and System 2 thinking patterns of human cognitive theory.
Preliminary and Observation
Background and Notations
LLMs.
@@
Given a prompted question , conventional LLM generates a response by iteratively predicting each response token conditioned on and generated tokens . This is termed as System 1 (Li et al. 2025b) thinking:
where [Prompt] and [Answer] are question and response , which is often short and concise.
LRMs.
Whereas in reasoning-focused LRMs, the generation process is split into two stages, e.g., thinking and conclusion, which is referred to as System 2 (Li et al. 2025b):
@@
where and are two delimiters denoting the start and end of reasoning, [Thoughts] and [Conclusion] are detailed reasoning process and concise answer, respectively. Formally, the reasoning part [Thoughts] reveals a trajectory of problem solving, known as CoT (Wei et al. 2022; Wang et al. 2023). At the micro level, some temporally consecutive tokens in [Thoughts] can be aggregated into short chunk based on semantics, logic and grammar:
@@
Similarly, each reasoning chunk in this paradigm is autoregressively generated by conditioning on question and previous reasoning thoughts: . This explicit decomposition not only improves interpretability through transparent exposure of reasoning steps but also enhances the model’s capability for complex tasks. Specifically, some chunks are well structured by prefix keywords such as “Wait”, “However”, “Alternatively” (Guo et al. 2025), marking the subsequent chunks will undergo logical backtracking, self-verification, or exploration. Works (Li et al. 2025a) have shown that reflection structures are crucial to LRMs. Under this setting, the “overthinking” phenomenon is that many seem delicate yet contribute nothing to reasoning logic, resulting in prominent overhead.
Observation: Reasoning Process Intervention
@@
To address the overthinking issue, a straightforward approach is difficulty reminders—encouraging LRMs to self-regulate output length based on task difficulty judgment through prompts like “Please think quickly if you find it easy”. However, this conventional interventional technique necessitates two prerequisites: accurate instruction following and robust task difficulty cognition. While prior studies (Wu et al. 2025; Han et al. 2025) have highlighted the inherent challenges of the former, we have already shown that current LRMs struggle to recognize varying levels of problem difficulty in Fig. 2.
@@
Instead of using a difficulty reminder in the prompt, we claim that a promising solution is transferring the intervention from the prompt to the inner reasoning process. We conduct a pre-experiment in Fig. 3, pretending the model has already correctly recognized the difficulty level of samples. Specifically, for a simple problem, we change the model input from “User: [Prompt]+[Question]. Assistant:” to “User: [Prompt]+[Question]. Assistant: This is a simple question, let’s think quickly.”. While for hard problems, the corresponding predefined response prefix is “It seems difficult, let’s think thoroughly.”. Fig. 3 shows that predefined difficulty awareness indeed helps to reduce response length, especially matched awareness (“simple, quickly” for easy and “difficult, thoroughly” for hard is matched, the reverse is unmatched).
@@
We define them as self-hypnosis that acts as a trigger in the output, guiding the subsequent inference trajectory. In fact, the reflection structures in LRMs (Guo et al. 2025; Li et al. 2025a) are essentially similar triggers aiming to intervene in subsequent think chunks. Compared to explicit prompts, we argue that these built-in triggers offer three key advantages: (1) native flexibility, (2) robustness, and (3) superior instruction-following capabilities. In the next section, we introduce how we guide LRMs to generate native interventional triggers of difficulty and redundancy cognition, autonomously.
Methodology
In this section, we describe our two-stage fine-tuning strategy Think-How-to-Think (TH2T) for mitigating overthinking, as shown in Fig. 4.
Stage 1: Difficulty Cognition Injection
@@
We aim to endow LRM with the capability to assess the difficulty of a given problem. We define two datasets: for easy problems, and for difficult ones. Two models from the same series were selected to ensure consistency in response: for basic LLM with prompted Short-CoT capability (e.g., Qwen2.5 series) and for advanced LRM (e.g., DeepSeek-R1-Distill-Qwen2.5 series). For each problem in , the responses are recorded as:
@@
where means the response of on is correct. can replicate the majority correct responses of on , which means: , wherein significant additional overhead yields limited performance gain. This motivates us to preserve ’s potential for efficiency on with satisfactory performance. Our desired thinking pattern of is responsed concisely on and elaborately on . To align with this objective, we construct a draft dataset grounded in difficulty differentiation:
where , similar to Eq. (5). aims to maintain the performance gain of on .
@@
To enable to perceive problem difficulty level autonomously, we introduce the trials of Fig. 3. We inject specific difficulty-aware cognitive Hypnosis into the prefix of answer within , which, as demonstrated before, functions as interventional triggers guiding the reasoning trajectory:
@@
If , the difficulty hypnosis H is defined as follows: “hypnosis This is a simple question, let’s think quickly./hypnosis”. If , then “hypnosisIt seems difficult, let’s think thoroughly./hypnosis”. The markers “hypnosis” and “/hypnosis” are defined as indicators for inserted difficulty hypnosis. The injected hypnosis is recognized as an embedded difficulty label to train to generate native difficulty cognition on the refined dataset , stimulating a concise thinking pattern on easy tasks. The execution of stage 1 is outlined at the upper section of the algorithmic framework Fig. (4).
Stage 2: Redundancy Cognition Injection

### v3

@@ Abstract
Recent Large Reasoning Models (LRMs) excel at complex reasoning tasks but often suffer from overthinking, generating overly long and redundant reasoning trajectories. To explore its essence, our empirical analysis reveals that LRMs are primarily limited to recognizing task properties (i.e., difficulty levels) like humans before solving the problem, leading to a one-size-fits-all reasoning strategy. This observation motivates a fundamental question: Can we explicitly bootstrap such ability to alleviate overthinking in LRMs? To this end, we propose Think-How-to-Think (TH2T), a novel two-stage fine-tuning strategy that progressively inspires LRMs’ difficulty cognition and redundancy cognition of LRMs. Specifically, we first inject Difficulty Dypnosis into output prefixes as cues for global, prospective reasoning strategy selection, stimulating the model’s sharper sensitivity to task complexity and adaptive control of reasoning depth. Then, we incorporate Redundancy Hypnosis into in-progress reasoning steps, which serve as local, retrospective signals for behavior correction by identifying and eliminating superfluous reasoning detours. Experiments across 7B/14B/32B models demonstrate that TH2T significantly reduces inference costs by over 70% on easy tasks and 40% on complex ones without compromising performance. The resultant models exhibit a nascent ability for difficulty-aware reasoning, effectively mitigating behaviors like excessive reflection and looping, thereby paving the way for more cognitively efficient LRMs.
Think How to Think: Mitigating Overthinking with Autonomous Difficulty Cognition in Large Reasoning Models
Yongjiang Liu1, Haoxi Li1, Xiaosong Ma2, Jie Zhang1, Song Guo1, 1The Hong Kong University of Science and Technology, 2The Hong Kong Polytechnic University
1 Introduction
@@
The emergence of Large Reasoning Models (LRMs) has led to substantial improvements in problem-solving capabilities Sun et al. (2024b); Yu et al. (2024); Team et al. (2025), particularly through the use of elaborate and sequential reasoning enabled by test-time scaling Snell et al. (2025). Many models, such as OpenAI-O1 OpenAI et al. (2024a), Deepseek-R1 Guo et al. (2025), QwQ Team (2025) leverage long Chain-of-Thought Wei et al. (2022); Wang et al. (2023) to approach complex tasks step-by-step, enabling a human-like thinking pattern with reflection, backtracking, and self-validation Xie et al. (2025); Liu et al. (2025). This reasoning paradigm demonstrably enhances the models’ problem-solving capabilities and yields promising results. However, lengthy responses introduce the problem of “overthinking” Chen et al. (2025), wherein model performance initially improves with extended reasoning chains but deteriorates beyond a certain length. This inefficiency is starkly evident when trivial queries, such as “What is 1 plus 1?”, trigger disproportionately complex reasoning paths. Such redundant deliberation not only inflates computational cost and latency but also degrades the user experience. These lead us to a central inquiry: Do LRMs possess the cognitive judgment to dynamically regulate their reasoning depth in response to varying problem difficulty?
@@
Inspiration for a solution lies in the human cognitive science: the dual-process theory of human reasoning. This theory posits a cognitive architecture composed of two distinct but complementary systems: a fast, concise, and intuitive System 1, and a slow, deliberate, and analytical System 2 Kahneman (2011); Evans and Stanovich (2013). Before engaging in problem-solving, humans often perform an initial assessment of a task’s difficulty, allowing them to allocate cognitive resources appropriately between these two systems Efklides (2011); Flavell (1979). For straightforward tasks, reliance on System 1 suffices, whereas complex problems necessitate the engagement of System 2. Our empirical analysis (Fig. 2) reveals that they approach disparate tasks—from the straightforward (e.g., GSM8K) to the highly complex (e.g., AIME2024)—with a monolithic and inefficient reasoning strategy. These limitations highlight the need for LRMs to develop autonomous difficulty cognition—to “think how to think”.
In this paper, we propose Think-How-to-Think (TH2T), a novel two-stage fine-tuning strategy that progressively inspires LRMs difficulty cognition and redundancy cognition. In contrast to prompt-based input guidance for instruction adherence, our approach emphasizes output-level interventions, directly intervening inside the reasoning trajectory to establish a self-hypnosis mechanism. In the first stage, we design Difficulty Hypnosis (DH) to stimulate the model’s awareness of task difficulty. DH introduces difficulty cues into output prefixes, which act as guiding signals for global, prospective strategy selection. With DH supervision from a hybrid dataset featuring both concise and elaborate reasoning trajectories, the model calibrates its sensitivity to task difficulty and adaptive control of reasoning depth, ultimately curtailing the impulse to overthink on straightforward problems through structured cognitive self-guidance.In the second stage, we design Redundancy Hypnosis (RH) to cultivate the model’s awareness of redundant reasoning. RH specifically targets fine-grained inference patterns such as redundant yet logically complete structures (e.g., reflections or double-checks) and terminal reasoning loops frequently observed in complex task generation. By embedding redundancy cues into these in-progress structures, RH provides local, retrospective signals for behavior correction, enabling the model to skip or truncate superfluous steps. In this way, the model strengthens its redundancy cognition and reduces superfluous reasoning detours. Overall, our contributions can be outlined as follows:
@@
We reveal that current LRMs fall short of perceiving task complexity. To mitigate overthinking, we propose enhancing LRMs with the ability to cognize and adapt to varying levels of problem difficulty autonomously.
@@
To align with human System 1 and System 2 thinking patterns, we propose a novel two-stage fine-tuning strategy with a self-hypnosis mechanism for prospective and retrospective intervention, thus progressively inspiring LRMs’ difficulty and redundancy cognition.
@@
Experiments demonstrate over 70% length reduction on easy tasks and 40% on hard tasks with maintained robust performance, as shown in Fig.1, exhibiting clear difficulty-aware capabilities and reduced redundancy.
2 Literature Review
@@
The advent of LRM OpenAI et al. (2024a); Guo et al. (2025); Team (2025) enables breakthroughs in problem solving Sun et al. (2024b); Yu et al. (2024) with Chain-of-Thought Wei et al. (2022); Wang et al. (2023) and test-time scaling Snell et al. (2025). Nevertheless, this introduces the challenge of overthinking Chen et al. (2025); Team et al. (2025); Sui et al. (2025).
@@
Prompt-based methods. Some works Aytes et al. (2025); Ding et al. (2024); Ning et al. (2024); Han et al. (2025); Xu et al. (2025); Lee et al. (2025) instruct LRMs by various prompts to respond concisely. Specifically, Renze and Guven (2024) explicitly prompts LLMs to perform concise step-by-step reasoning. SoT Ning et al. (2024) combines reasoning paradigms with linguistic constraints to minimize token usage. Token-Budget Han et al. (2025) establishes a self-estimated token budget as a constraint. Ge et al. (2025) regulates thinking token distribution by few-shot in-context learning.
Output-based methods. Implicit-based methods Deng et al. (2024); Hao et al. (2024); Saunshi et al. (2025); Cheng and Durme (2024); Shen et al. (2025a) propose to internalize reasoning steps within the latent representation, thereby shifting the cognitive process from the explicit language space to the implicit hidden space. Saunshi et al. (2025) simulates multiple steps of CoT via a looped transformer layer to deepen the reasoning process internally. Another line of work focuses on dynamic execution. Sun et al. (2024a); Xie et al. (2023); Liao et al. (2025) optimize Best-of-N (BoN) decoding with speculation and rejection. Li et al. (2024); Yang et al. (2025a); Wang et al. (2025) explores potential early exit point. NoThinking Ma et al. (2025a) prompts reasoning models to directly output final answers without thinking.
Model-based methods. Fine-tuning offers a direct post-training strategy Yang et al. (2025b). CoT-Value Ma et al. (2025b) learns a controllable direction in the parameter space to steer generation length. C3oT Kang et al. (2025) employs GPT-4 as a compressor while retaining key information. TokenSkip Xia et al. (2025) fine-tunes the model to prune non-essential tokens during inference. Alternatively, reinforcement learning provides another vulnerable and expensive paradigm, often by incorporating length penalties to balance performance and efficiency Arora and Zanette (2025); Team et al. (2025); Aggarwal and Welleck (2025). O1-pruner Luo et al. (2025b) integrates a Length-Harmonizing Reward into a PPO-style loss. DAST Shen et al. (2025b) employs SimPO Meng and Xia (2024) on a constructed length-preference dataset, while AdaR1 Luo et al. (2025a) optimizes a merged hybrid long/short CoT model using DPO Rafailov et al. (2023).
Limitations and our approach. Overall, existing approaches to efficient reasoning largely operate under a monolithic paradigm of length compression, which introduces critical limitations: 1) Prompt-based methods exhibit unreliability of instruction following and task cognition; 2) Implicit output-based methods suffer from performance degradation and compromised interpretability; 3) Model-based approaches are constrained by a behavioral rigidity that hinders adaptability. In contrast, we reframe overthinking not just as a behavioral flaw, but as a fundamental cognitive deficit, since length is not an accurate proxy for difficulty. Our approach instills a dual-level awareness: a global, prospective cognition of task difficulty to guide high-level strategy, and a local, retrospective in-progress perception of redundancy for fine-grained execution. This self-aware intervention mechanism cultivates the metacognitive skill of thinking about how to think efficiently, rather than simply enforcing brevity.
3 Preliminary and Observation
3.1 Background
@@
LLMs. Given a prompted question , LLM generates a response by iteratively predicting each response token conditioned on and generated tokens . This is termed as System 1 thinking Li et al. (2025b): , where [Prompt] and [Answer] are question and response , which is often short and concise.
@@
LRMs. Whereas in LRMs, the generation process is split into two stages, i.e., thinking and conclusion, referred to as System 2: , where and are two delimiters denoting the start and end of reasoning, [Thoughts] and [Conclusion] are detailed reasoning processes and concise answers, respectively. Formally, [Thoughts] reveals a trajectory of problem solving, known as CoT Wei et al. (2022); Wang et al. (2023). At the micro level, some consecutive tokens in [Thoughts] can be aggregated into short chunk based on semantics, logic and grammar:
@@
Similarly, each reasoning chunk is autoregressively generated by conditioning on question and previous chunks: . This explicit decomposition not only improves interpretability through transparent exposure of reasoning steps but also enhances the model’s capability for complex tasks. Specifically, some reflective chunks are well structured by prefix keywords such as “Wait”, “However”, “Alternatively” Guo et al. (2025), marking the subsequent chunks will undergo logical backtracking, self-verification, or exploration, which have been proved crucial to LRMs Li et al. (2025a). Under this setting, we define “overthinking” as a phenomenon characterized by inefficiencies at both the global and local levels: globally, through an excessive number of reasoning steps () relative to task complexity, and locally, where individual steps () are elaborate yet logically superfluous, resulting in prominent overhead.
3.2 Reasoning Process Intervention
@@
To address the overthinking issue, a straightforward approach is difficulty reminders—encouraging LRMs to self-regulate output length based on task difficulty judgment through prompts like “Please think quickly if you find it easy”. However, this hypothetical condition necessitates two prerequisites: accurate instruction following and robust task difficulty cognition. While prior studies Wu et al. (2025); Han et al. (2025) have highlighted the inherent challenges of the former, we have already shown that current LRMs lack cognition of problem difficulty in Fig. 2.
@@
Instead, we reveal that a promising solution is transferring the intervention from the prompt to the inner reasoning process. We conduct a pre-experiment in Fig. 3, pretending the model knows the difficulty level of samples. Given a simple problem, we change the model input from “User: [Prompt]+[Question]. Assistant:” to “User: [Prompt]+[Question]. Assistant: This is a simple question, let’s think quickly.”. While for hard problems, the predefined response prefix is “It seems difficult, let’s think thoroughly.”. Fig. 3 shows that predefined difficulty awareness largely outperforms reminder on reducing token length, especially matched awareness (“simple, quickly” for easy and “difficult, thoroughly” for hard is matched, the reverse is unmatched).
@@
We define them as self-hypnosis that acts as triggers in the output, guiding the subsequent inference trajectory. Compared to explicit prompts, these built-in triggers offer three key advantages: 1) native flexibility; 2) robustness; 3) superior instruction-following capabilities. In Sec.4, we utilize the self-hypnosis mechanism to guide LRMs to generate native interventional triggers of difficulty and redundancy cognition, autonomously.
4 Methodology
As discussed in Sec.3.1, the challenge of overthinking stems from failures in both high-level strategy selection and fine-grained step execution. We ground our method in the dual-process theory of cognition, i.e., intuitive System 1 and analytical System 2. The strategy calls for a global, prospective intervention applied pre-generation, while the execution demands a local, retrospective mechanism for in-progress correction and refinement. Since they are complementary and distinct systems, we propose a two-stage fine-tuning framework Think-How-to-Think (TH2T) for mitigating overthinking, as shown in Fig. 4.
4.1 Stage 1: Difficulty Cognition Injection
@@
The foundational principle of our approach is to cultivate the model’s capacity for task appraisal, which is a metacognitive prerequisite for adaptive reasoning. We define two datasets: for easy problems, and for difficult ones. Two models from the same series were selected to ensure consistency in response: for basic LLM with prompted Short-CoT capability and for advanced LRM. For each problem in , the responses are recorded as:
@@
where means the response of on is correct. can replicate the majority correct responses of on , which means: , wherein significant additional overhead yields limited performance gain. This motivates us to preserve ’s potential for efficiency on with satisfactory performance. The target reasoning pattern of entails concise responses on and detailed responses on . To align with this objective, we construct a draft dataset grounded in difficulty differentiation:
where . aims to maintain the performance gain of on .
@@
To enable to perceive problem difficulty level autonomously, we introduce the trials of Sec. 3.2. We inject specific difficulty-aware cognitive Hypnosis into the prefix of answer within , which, as demonstrated before, act as an interventional signals for global, prospective strategy selection.
@@
If , the Difficulty Hypnosis H is defined as follows: “hypnosis This is a simple question, let’s think quickly./hypnosis”. If , then “hypnosisIt seems difficult, let’s think thoroughly./hypnosis”. The markers “hypnosis” and “/hypnosis” are indicators for inserted difficulty hypnosis. The injected hypnosis is recognized as an embedded difficulty label to train to generate native difficulty cognition, stimulating a concise thinking pattern on easy tasks. The execution of stage 1 is outlined in the upper section of Fig. 4.
4.2 Stage 2: Redundancy Cognition Injection